OpenStudy (anonymous):

solve : 1) (x+y^(2) * x) dx + e^(x^e) y dy=0 2) y-(2x-4y^2)dy/dx=0

OpenStudy (turingtest):

\[(x+y^2x)dx+(e^{x^e}y)dy=0\] is the first?

OpenStudy (anonymous):

yes

OpenStudy (mathteacher1729):

Separable... no. Integrating factor... not any obvious ones... euler equations... no. Gauss's equations.. no. My goodness. I honestly don't know, this looks to be something beyond a first year ordinary linear differential equation. What class is this for?

OpenStudy (turingtest):

Yeah I got nothing...

OpenStudy (mathteacher1729):

They're both nonlinear. I guess you can do numerical approximations on them if you have a decent idea of what might be a solution. I can't think of any physical situations where such differential equations might arise. I'm curious to know what class these equations come from. http://math.exeter.edu/rparris/winplot.html is a great freeware program which should let you qualitatively analyze them (graph them, see specific solutions, etc.). It also has a good numerical solver. Sorry I can't be much help beyond that.

OpenStudy (jamesj):

The first equation is seriously ugly and doesn't have a solution in terms of elementary functions; although you can force it look like this \[2yy' - 2xe^{x^e}y^2 = -2xe^{x^e}\] The left hand side can now be made exact by multiplying through by \[H(x) = \exp \left[ \int\limits -2xe^{x^e} dx \right]\] so we have \[\left(y^2 H(x) ) \right)' = -2xe^{x^e}H(x)\] and you can take it from there. With a little manipulation you can make H(x) look like something involving the Gamma Function, but it is well known that function is not reducible to elementary functions.

OpenStudy (jamesj):

The second equation is slightly easier but I'm also convinced it doesn't have a solution in terms of elementary functions. I thought of this second equation first as an equation in x = x(y) and dx/dy. Then it became relatively easy to solve for x. Inverting that resulting equation x=x(y) is unfortunately not straight-forward.

OpenStudy (turingtest):

Dang James you are the man. Wow...

OpenStudy (anonymous):

thanks , i will see wat i will get but as i got this equation need to do this equation through linear due to y or x or Bernoulli or Separable one of these ways :S

OpenStudy (mathteacher1729):

Respect to James. :) Nicely done!

OpenStudy (jamesj):

To confirm also, the second equation is \[{dy \over dx} = {y \over 2x-4y^2}\] yes?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

I was hoping it would be a Bernoulli equation for a moment, but you can there's no obvious transformation to put this in Bernoulli form.

OpenStudy (anonymous):

then i guess we make it dx/dy = (2x-4y^2 )/ y

OpenStudy (jamesj):

Right, that's what I did and now it's straightforward

OpenStudy (anonymous):

yes :)